OpenStudy (anonymous):

what is 2 over 5 + 7 over 8 ?

OpenStudy (anonymous):

2/5+7/8= 51/40 51/40=1(11/40)

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

\[51/40 = 1_{40}^{11} :)\]